OpenStudy (anonymous):

Find the next term of the sequence: 1/128, -1/16, 1/2, -4, 32

OpenStudy (anonymous):

Find the common ratio first..

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

well, because he 1 is in front and the 4 is in the middle of the sequence?

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Common ratio is the ratio between two successive terms, this is common which means equal..

OpenStudy (anonymous):

your first term is : 1/128 your second terms is : -1/16

OpenStudy (anonymous):

\[Common \; Ratio = \frac{\frac{-1}{16}}{\frac{1}{128}} = \frac{-128}{16} = -8\]

OpenStudy (anonymous):

So than this would be the correct answer right -256

OpenStudy (anonymous):

Yes it is..